# Accelerating Materials Design via LLM-Guided Evolutionary Search

### v1

@@ 2.5 Implementation Details
LLEMA is designed for seamless evaluation on new tasks and minimal engineering overhead. To evaluate a new task, the user is only required to supply a CSV file specifying the task name and associated property constraints. From this input, the agentic framework automatically and iteratively constructs prompts to generate candidate materials, followed by their crystallographic structures in CIF format using an LLM. Next, for property evaluation, LLEMA adopts a hierarchical oracle strategy where the candidates are first queried against curated materials databases, while pretrained ML surrogates (e.g., CGCNN, ALIGNN) are invoked only for out-of-distribution compounds. These surrogates are used strictly in inference mode with publicly available pretrained weights, which avoids retraining and significantly reduces computational overhead. For reproducibility, we report all surrogate model checkpoints and APIs in Table 4 (Appendix D.2). Candidates violating hard constraints are directly assigned low scores, ensuring efficient pruning before expensive evaluations. This design enables scaling across tasks and large candidate spaces. A detailed implementation of LLEMA is provided in Appendix D.1.
@@ 3.1 Datasets and Benchmarks
We evaluate LLEMA on fourteen application-driven discovery tasks spanning electronics, energy, coatings, optics, and aerospace, to probe multi-objective reasoning under realistic constraints, and thermodynamic stability (Table 1). Each task reflects an industrially relevant challenge and is designed to probe multi-objective reasoning under realistic constraints. Our benchmark design follows three guiding principles: (i) Application relevance: the target properties align with pressing technological needs such as sustainable energy, advanced electronics, and structural resilience; (ii) Multi-constraint optimization: tasks involve simultaneous optimization of multiple, often competing, objectives (e.g., maximizing hardness and conductivity), mirroring real-world engineering specifications; and (iii) Thermodynamic stability: all tasks enforce stability requirements through formation energy and energy-above-hull criteria to ensure synthetic accessibility rather than purely theoretical feasibility. Appendix B contains additional details on all tasks, including property thresholds and predictive models.
Table 1: Material Discovery Benchmark. Each task is characterized by its application domain and quantitative property constraints.
@@
Wide-Bandgap Electronics Band gap $\geq 2.5$ eV; Formation energy $\leq-1.0$ eV/atom
Semiconductors Energy above hull $\leq 0.1$ eV/atom
@@
Energy above hull $\leq 0.1$ eV/atom; Must contain Li, Na, K, Mg, Ca, or Al
Piezo Energy Harvesters Energy Piezoelectric coefficient $\geq 8$ pC/N ; Dielectric constant $10\leq\kappa\leq 8000$
@@
Photovoltaic Absorbers
Energy
Band gap 0.70.72.02.0 eV; Formation energy $\leq 0.0$ eV/atom
Composed of earth-abundant elements only
Hard Coating Materials Mechanical Bulk modulus 200200500500 GPa; Shear modulus 100100300300 GPa
@@
Structural Materials Aerospace Bulk modulus 100100300300 GPa; Shear modulus 6060200200 GPa
for Aerospace Formation energy <0<0 eV/atom
Acousto-Optic Hybrids
Acoustics / Optics
Piezoelectric coefficient: $2$ - $9$ pC/N ; Dielectric constant: 88-$95$
@@
Perovskite Oxides Sustainability Exclude Pb, Cd, Hg, Tl, Be, As, Sb, Se, U, Th; Prefer stable ABO3 oxide structures
@@ 3.2 Experimental Setup
We assessed LLEMA using complementary evaluation criteria designed to capture both efficiency and quality of material generation. Hit-Rate measures the percentage of generated candidates that simultaneously satisfy all property constraints, quantifying the efficiency of valid discovery. Stability evaluates the percentage of valid and thermodynamically stable, reflecting physical practicality beyond theoretical feasibility. Specifically, materials having energy above hull value less than $0.1$ eV/atom are considered stable. Finally, Pareto Front Analysis compares the quality of multi-objective trade-offs, with superior methods producing non-dominated solutions that span larger and more diverse regions of the design space. We benchmark LLEMA against generative models such as CDVAE (Xie et al., 2021), G-SchNet (Gebauer et al., 2019), and DiffCSP (Jiao et al., 2023), as well as LLM-driven approaches including LLMatDesign (Jia et al., 2024) and direct prompting baselines. To ensure budget fairness, all non-LLM baselines (e.g., CDVAE, G-SchNet, DiffCSP) were allotted 10× more candidate generations ($15,000$ vs. $1,500$) for LLM based methods to offset their lack of in-context feedback and higher computational cost per sample. Unlike the baselines, LLEMA refines its outputs by iteratively sampling candidates from its experience buffer through an in-context refinement process. Detailed implementation settings for baselines are provided in Appendix C.
@@ 3.3 Quantitative Results
Table 2: Comparison of Baselines on Materials Discover Benchmark. We implemented with GPT-4o-mini and Mistral-Small-3.2-24B-Instruct-2506, against state-of-the-art baselines across materials design tasks. We report Hit Rate (H.R.) and Stability (Stab.), where higher values indicate better performance.
Method
Wide-Bandgap
Semicond.
SAW/BAW
Acoustic Substrates
High-$k$
Dielectrics
Solid-State
Electrolytes
Piezo Energy
Harvesters
Transparent
Conductors
Insulating
Dielectrics
@@
Method
Photovoltaics
Absorbers
Hard Coatings
Materials
Hard, Stiff
Ceramics
Aerospace
Materials
Acousto-optic
Hybrids
Low Density
Structures
Perovskite
Oxides
@@
LLEMA (GPT) 22.90 4.76 17.78 4.61 60.99 5.73 0.97 0.26 26.26 0.82 0.47 0.14 19.37 2.79
Table 2 reports task-specific performance across fourteen benchmark domains, measured by hit rate (H.R) and stability (Stab.) under varying physical and chemical constraints. Overall, LLEMA consistently outperforms all baselines, achieving both higher hit rates and markedly better stability across diverse material classes. Traditional generative models perform reasonably on simpler tasks like wide-bandgap semiconductors or photovoltaic absorbers but often produce valid yet unstable candidates. Even LLM-based baselines show limited robustness, performing well in some domains but failing in others, suggesting poor generalization across different physical regimes. In contrast, LLEMA consistently achieves the highest hit rate and markedly higher stability compared to the baselines. This elevated stability indicates that evolutionary refinement and chemistry-informed rules enable LLEMA to not only meet design constraints more reliably but also generate thermodynamically consistent, physically meaningful structures. Overall, these results demonstrate that LLEMA generalizes robustly across material classes, combining exploration and constraint satisfaction more effectively than prior methods. A deeper discussion on the qualitative aspects of LLEMA follows in the upcoming sections.
@@ 3.4 Qualitative Results
We analyzed qualitative outcomes to understand LLEMA’s behavior under realistic discovery settings, focusing on: (i) Convergence dynamics, which examine how iterative feedback progressively steers the LLM toward feasible design regions; (ii) Pareto trade-offs, which assess whether the method can balance competing objectives under strict property constraints; and (iii) Discovered candidates, which illustrate how novel yet chemically plausible compositions emerge and how they align with domain knowledge.
@@
Figure 3: Evolution of candidates for the Stable Wide-Bandgap Semiconductor at different stages of LLEMA.
First, we analyzed how solutions evolve over iterations. Early generations scatter broadly across the search space, often violating property constraints or collapsing into suboptimal regions. As the search progresses, feedback from the oracle, evolutionary memory, and constraint-based selection gradually steer LLEMA toward more promising areas of the chemical design space. The proportion of valid candidates increases from roughly 27% at the 250th iteration to about 33% near the 1000th iteration, indicating that the algorithm effectively exploits high-fitness regions while maintaining exploration across diverse chemical configurations. This gradual improvement reflects stronger selection pressure toward feasible trade-offs, resulting in a visible migration of candidate clusters into constraint-satisfying zones (e.g., band gaps above $2.5$ eV with low formation energies). These convergence dynamics show how iterative, feedback-driven evolution refines the LLM’s proposal distribution, preserving diversity while increasingly focusing on candidates that balance multiple objectives in physically meaningful ways.
@@
Figure 4: Pareto front analysis of candidate materials for two design tasks. (a) Wide-Bandgap Semiconductors; (b) Hard–Stiff Ceramics.
We then examined the Pareto fronts for the Wide-Bandgap Semiconductor and Hard–Stiff Ceramic tasks, both of which impose stringent property requirements: semiconductors must exhibit band gaps $\geq 2.5$ eV with formation energies $\leq-1.0$ eV/atom to ensure functionality and stability, while ceramics require bulk moduli in the range 100100300300 GPa and shear moduli between 6060200200 GPa. As shown in Figure 4, the optimal Pareto front is entirely dominated by LLEMA, with all Pareto-optimal solutions originating from our method across both tasks. This demonstrates that LLEMA not only generates a higher proportion of valid and thermodynamically stable candidates but also consistently identifies the globally optimal trade-offs between competing objectives. By enforcing explicit constraints and applying domain-guided evolutionary refinement, LLEMA effectively filters infeasible solutions while converging toward the true Pareto-optimal frontier, surpassing all baseline methods that fail to reach this region of the design space.
@@
Finally, we evaluated the plausibility of discovered compositions in the real world. We find that materials proposed by LLEMA align with families previously investigated by domain experts, underscoring its ability to generate realistic candidates. For instance, in the High-$k$ dielectrics task, LLEMA suggests ZrAl2O5 and Hf0.5Zr0.5O2, which connect closely to Zr–Al and Hf–Zr oxides studied as promising high-$k$ materials (Hakala et al., 2006; Das & Jeon, 2020; Islam et al., 2021). Similarly, LLEMA reflects expert strategies such as substitution and doping, e.g., proposing BaHfZr oxide, consistent with known dopant-driven improvements in HfZr oxides (Kim et al., 2024). In photovoltaics, candidates such as CaZnSi and MgZnSi oxides emerge, which, while not directly reported, are chemically related to established ZnO-based systems (Esgin et al., 2022). These examples demonstrate that LLEMA not only respects constraints but also uncovers novel yet chemically plausible families, validating its utility for guiding real-world discovery.
@@ 4 Analysis
4.1 Memorization vs. Guided Exploration
A central challenge in leveraging large language models (LLMs) for scientific discovery is their tendency to memorize training data and regenerate it during generation, rather than exploring novel solutions. Prior work has shown that LLMs frequently reproduce examples from their training corpus (Carlini et al., 2021; Hartmann et al., 2023). In materials discovery, this manifests as repeated suggestions of compounds already present in databases such as the Materials Project, leading to high duplication rates and limited novelty. We compare three approaches: a direct LLM call, an LLM augmented with a memory buffer and iterative feedback, and LLEMA. The direct LLM call exhibits the highest duplication and near-total reliance on the Materials Project (e.g., High-$k$ dielectrics show almost 100% overlap). While adding a memory buffer helps the model explore diverse leads, it still shows a high rate of memorization, thus implying that simply storing past candidates does not guide the search toward new or diverse regions of chemical space. In contrast, incorporating a multi-island evolutionary framework enables the model to avoid local optima and repeated patterns. When further combined with chemically informed rules such as oxidation-state consistency, stoichiometry preservation, and prototype substitution, LLEMA effectively reduces redundancy and expands exploration into novel, chemically plausible regions of the design space. Together, these components push the model beyond corpus recall toward genuine discovery.
@@ 4.2 Impact of Domain-Guided Evolutionary Refinement
Table 3: Comparison of hit-rate (H.R), stability (Stab.) and Memorization Rate (Mem.) across generation methods aggregated over four datasets.
@@
We evaluate the benefits of evolutionary refinement with domain-guided generation rules by benchmarking LLEMA against two baselines: (i) an LLM with iterative feedback (LLM w/ Memory), and (ii) an unguided mutation–crossover search implemented within a multi-island evolutionary framework following Romera-Paredes et al. (2024). All methods were run for $250$ iterations across four benchmark datasets, with property constraints relaxed by 20% to allow broader exploration while preserving task relevance. Results from Table 3 show that LLMs with iterative feedback yield limited improvement, as the model tends to recall known materials and overfit to training patterns. However, the introduction of multi-island evolution substantially improves hit rate and stability by promoting parallel exploration and mitigating premature convergence, though the absence of chemical constraints results in memorization. Incorporating chemistry-informed generation rules in LLEMA achieves the best overall balance (H.R = $30.2$, Stab. = $27.6$, Mem. = $16.6$), constraining the search to thermodynamically and compositionally plausible regions while maintaining diversity. This staged refinement from using memory to evolutionary search and domain-guided evolution demonstrates how each component progressively enhances exploration, stability, and chemical realism in generative materials design.

### v2

@@ 2.5 Implementation Details
To evaluate a new task, the user is only required to supply a CSV file specifying the task name and associated property constraints. From this input, the agentic framework automatically and iteratively constructs prompts to generate candidate materials, followed by their crystallographic structures in CIF format using an LLM. Next, for property evaluation, LLEMA adopts a hierarchical oracle strategy where the candidates are first queried against curated materials databases, while pretrained ML surrogates (e.g., CGCNN, ALIGNN) are invoked only for out-of-distribution compounds. These surrogates are used in inference mode with publicly available pretrained weights, which avoids retraining and significantly reduces computational overhead. For reproducibility, we report all surrogate model checkpoints and APIs in Table 5 (Appendix B.1). Candidates violating hard constraints are directly assigned low scores, ensuring efficient pruning before expensive evaluations. A detailed implementation of LLEMA is provided in Appendix D.2.
@@ 3.1 Datasets and Benchmarks
We evaluate LLEMA on fourteen application-driven discovery tasks spanning electronics, energy, coatings, optics, and aerospace, to probe multi-objective reasoning under realistic constraints, and thermodynamic stability (Table 2). Each task reflects an industrially relevant challenge and is designed to probe multi-objective reasoning under realistic constraints. Our benchmark design follows three guiding principles: (i) Application relevance: the target properties align with pressing technological needs such as sustainable energy, advanced electronics, and structural resilience; (ii) Multi-constraint optimization: tasks involve simultaneous optimization of multiple, often competing, objectives (e.g., maximizing hardness and conductivity), mirroring real-world engineering specifications; and (iii) Thermodynamic stability: all tasks enforce stability requirements through energy-above-hull criteria ($E_{hull}\approx 0$) to ensure synthetic accessibility rather than purely theoretical feasibility. Appendix C contains additional details on all tasks, including property thresholds and predictive models.
Table 2: Materials Discovery Benchmark. Each task is characterized by its application domain and quantitative property constraints.
@@
Wide-Bandgap Electronics Band gap $\geq 2.5$ eV
Semiconductors Formation energy $\leq-1.0$ eV/atom
@@
Must contain Li, Na, K, Mg, Ca, or Al
Piezo Energy Harvesters Energy Piezoelectric coefficient $\geq 8$ pC/N; $10\leq\kappa\leq 8000$
(If ranking: use FoM $d^{2}/\kappa$; higher is better)
@@
Photovoltaic Absorbers Energy Band gap 0.70.72.02.0 eV; Formation energy $\leq 0.0$ eV/atom
Earth-abundant and non-toxic elements only
Hard Coating Materials Mechanical Bulk modulus $\geq 200$ GPa; Formation energy $\leq-1.0$ eV/atom; Band gap $\geq 3.0$ eV
@@
Structural Materials Aerospace Density $\leq 5.0$ g/cm3; Bulk modulus $\geq 100$ GPa
for Aerospace Shear modulus $\geq 40$ GPa; Energy above hull $\leq 5.0$ eV/atom
Acousto-Optic Hybrids Acoustics / Optics $3\leq$ Piezoelectric coefficient $\leq 9$ pC/N; $8\leq\kappa\leq 85$
@@
Perovskite Oxides Sustainability Exclude Pb, Cd, Hg, Tl, Be, As, Sb, Se, U, Th; Prefer stable ABO3 oxides
@@ 3.2 Experimental Setup
We assessed LLEMA using complementary evaluation criteria designed to capture both the efficiency and the quality of material generation. Hit-Rate measures the percentage of generated candidates that simultaneously satisfy all property constraints, quantifying the efficiency of valid discovery. Stability evaluates the percentage of valid and thermodynamically stable, reflecting physical practicality beyond theoretical feasibility. Specifically, materials having an energy above the hull value less than $0.1$   eV/atom are considered stable. Finally, Pareto Front Analysis compares the quality of multi-objective trade-offs, with superior methods producing non-dominated solutions that span larger and more diverse regions of the design space. We benchmark LLEMA against generative models such as CDVAE (Xie et al., 2022), G-SchNet (Gebauer et al., 2019), DiffCSP (Jiao et al., 2023), and MatterGen (Zeni et al., 2025), as well as LLM-driven approaches including LLMatDesign (Jia et al., 2024) and direct prompting baselines. For comparison, all non-LLM baselines (e.g., CDVAE, G-SchNet, DiffCSP) were assigned $\sim$ 10× candidate generations than LLM-based methods, generating 1500 samples, to offset their lack of in-context feedback. Unlike the baselines, LLEMA refines its outputs by iteratively sampling candidates from its experience buffer through an in-context refinement process. Detailed implementation settings for baselines are provided in Appendix D.
@@ 3.3 Quantitative Results
Table 3 reports task-specific performance across fourteen benchmark domains, measured by hit-rate (H.R) and stability (Stab.) under varying physical and chemical constraints. Overall, LLEMA consistently outperforms all baselines, achieving higher hit-rates and markedly better stability across diverse material classes. Traditional generative models perform reasonably well on tasks like ‘hard, stiff ceramics’ or ‘acousto-optic hybrids’ but often produce valid yet unstable candidates. Even LLM-based baselines show limited robustness, performing well in some domains but failing in others, suggesting poor generalization across different physical regimes. In contrast, LLEMA consistently achieves higher hit-rates and markedly greater stability than the baselines. This elevated stability indicates that evolutionary refinement and chemistry-informed rules enable LLEMA not only to meet design constraints more reliably but also to generate thermodynamically consistent and physically meaningful structures. Overall, these results demonstrate that LLEMA generalizes robustly across material classes, combining exploration and constraint satisfaction more effectively than prior methods. A deeper discussion on the qualitative aspects of LLEMA follows in the upcoming sections.
Table 3: Comparison of Baselines on Materials Discovery Benchmark. We implemented LLEMA with GPT-4o-mini and Mistral-Small-3.2-24B-Instruct-2506, against state-of-the-art baselines across materials design tasks. We report hit-rate (H.R.) and stability (Stab.), where higher values indicate better performance.
Method Wide-Bandgap Semicond. SAW/BAW Acoustic Substrates High-$k$ Dielectrics Solid-State Electrolytes Piezo Energy Harvesters Transparent Conductors Insulating Dielectrics
@@
Method Photovoltaics Absorbers Hard Coating Materials Hard, Stiff Ceramics Aerospace Materials Acousto-optic Hybrids Low Density Structures Perovskite Oxides
@@
LLEMA (GPT) 22.90 4.76 17.78 4.61 30.99 5.73 0.97 0.26 26.26 0.82 0.47 0.14 19.37 2.79
@@ 3.4 Qualitative Results
We analyzed qualitative outcomes to understand LLEMA’s behavior under realistic discovery settings, focusing on: (i) Convergence dynamics, which examines how iterative feedback progressively steers the LLM toward feasible design regions; (ii) Pareto trade-offs, which assess whether the method can balance competing objectives under strict property constraints; and (iii) Discovered candidates, which illustrate how novel yet chemically plausible compositions emerge and how they align with domain knowledge.
@@
Figure 2: Evolution of candidates for the Wide-Bandgap Semiconductor task at different stages of LLEMA.
First, we analyzed how solutions evolve over iterations. Early generations scatter broadly across the search space, often violating property constraints or collapsing into suboptimal regions. As the search progresses, the feedback from the oracle, evolutionary memory, and constraint-based selection gradually steers LLEMA toward more promising areas of chemical design space. The proportion of valid candidates increases from roughly 27% at the 250th iteration to about 33% near the 1000th iteration, indicating that the algorithm effectively exploits high-fitness regions while maintaining exploration across diverse chemical configurations. This gradual improvement reflects stronger selection pressure toward feasible trade-offs, resulting in a visible migration of candidate clusters into constraint-satisfying zones (e.g., band gaps above $2.5$ eV with low formation energies). These convergence dynamics show how iterative feedback-driven evolution refines the LLM’s proposal distribution, preserving diversity while increasingly focusing on candidates that balance multiple objectives in physically meaningful ways.
@@
Figure 3: Pareto front analysis of candidate materials for two design tasks. (a) Wide-Bandgap Semiconductors; (b) Hard, Stiff Ceramics.
We then examined the Pareto fronts for the Wide-Bandgap Semiconductor and Hard–Stiff Ceramic tasks, both of which impose stringent property requirements: semiconductors must exhibit band gaps $\geq 2.5$ eV with formation energies $\leq-1.0$ eV/atom to ensure functionality and stability, while ceramics require bulk moduli in the range 100100300300 GPa and shear moduli between 6060200200 GPa. As shown in Figure 3, the optimal Pareto front is completely dominated by LLEMA, and all Pareto-optimal solutions originate from our method on both tasks. This shows that LLEMA not only generates a higher proportion of valid and thermodynamically stable candidates, but also consistently identifies the globally optimal trade-offs between competing objectives. By enforcing explicit constraints and applying domain-guided evolutionary refinement, LLEMA effectively filters infeasible solutions while converging toward the true Pareto-optimal frontier, surpassing all baseline methods that fail to reach this region of the design space.
@@
Finally, we evaluated the plausibility of discovered compositions in the real world. We find that the materials proposed by LLEMA align with families previously investigated by domain experts, underscoring their ability to generate realistic candidates. For example, in the High-$k$ dielectric task, LLEMA suggests ZrAl2O5 and Hf0.5Zr0.5O2, which connect closely to Zr–Al and Hf–Zr oxides studied as promising high-$k$ materials (Hakala et al., 2006; Das and Jeon, 2020; Islam et al., 2021). Similarly, LLEMA reflects expert strategies, such as substitution and doping, e.g., proposing BaHfZr oxide, which is consistent with known dopant-driven improvements in HfZr oxides (Kim et al., 2024). In photovoltaics, candidates such as CaZnSi and MgZnSi oxides emerge, which, although not directly reported, are chemically related to established ZnO-based systems (Esgin et al., 2022). These examples demonstrate that LLEMA not only respects constraints, but also uncovers novel yet chemically plausible families, validating its utility to guide real-world discovery.
@@ 4.1 Impact of Domain-Guided Evolutionary Refinement
Table 4: Comparison of hit-rate (H.R), stability (Stab.), and Memorization Rate (Mem.) across generation methods aggregated over four tasks.
@@
We evaluate the benefits of evolutionary refinement with domain-guided generation rules by benchmarking LLEMA against two baselines: (i) an LLM with iterative feedback (LLM w/ Memory), and (ii) an unguided mutation–crossover search implemented within a multi-island evolutionary framework following Romera-Paredes et al. (2024). All methods were run for $250$ iterations across four benchmark datasets, with property constraints relaxed by 20% to allow broader exploration while preserving task relevance. The results of Table 4 show that LLMs with iterative feedback yield limited improvement, as the model tends to recall known materials and overfit to training patterns. However, the introduction of multi-island evolution substantially improves hit-rate and stability by promoting parallel exploration and mitigating premature convergence, though the absence of chemical constraints results in memorization. Incorporating chemistry-informed generation rules in LLEMA achieves the best overall balance (H.R = $30.2$, Stab. = $27.6$, Mem. = $16.6$), constraining the search to thermodynamically and compositionally plausible regions while maintaining diversity. This staged refinement from the use of memory to the evolutionary search and domain-guided evolution demonstrates how each component progressively enhances exploration, stability, and chemical realism in generative materials design.
4.2 Memorization vs. Guided Exploration
A central challenge in leveraging large language models (LLMs) for scientific discovery is their tendency to memorize training data and regenerate it during generation, rather than exploring novel solutions. Prior work has shown that LLMs frequently reproduce examples from their training corpus (Carlini et al., 2021; Hartmann et al., 2023). In materials discovery, this manifests as repeated suggestions of compounds already present in databases such as the Materials Project, leading to high duplication rates and limited novelty. We compare three approaches: a direct LLM call, an LLM augmented with a memory buffer and iterative feedback, and LLEMA. The direct LLM call exhibits the highest duplication and near-total reliance on the Materials Project (e.g., High-$k$ dielectrics show almost 100% overlap). Although adding a memory buffer helps the model explore diverse leads, it still shows a high rate of memorization, thus implying that simply storing past candidates does not guide the search toward new or diverse regions of chemical space. In contrast, incorporating a multi-island evolutionary framework enables the model to avoid local optima and repeated patterns. When further combined with chemically informed rules such as oxidation-state consistency, stoichiometry preservation, and prototype substitution, LLEMA effectively reduces redundancy and expands exploration into novel, chemically plausible regions of the design space. Together, these components push the model beyond memorization toward genuine discovery.